Calvin invested $7500 for one year, part at 12% annual interest and the rest at 10% annual interest. his total interest fot the year was $890. How much money did he invest at 12%?
Answer by checkley77(6671)

.12x+.10(7,500-x)=890
.12x+750-.10x=890
.02x=890-750
.02x=140
x=140/.02
x=$7,000 is the amount invested @ 12%.